You burn 20 calories per minute biking for x minutes and 10 calories per minute walking for y minutes. You spend a total of 90 minutes biking and walking and burn 1300 calories.
Write a system of equations to determine how much time you spend on each exercise.
How many minutes did you spend biking?

Answer by macston(5194)

.
x+y=90
y=90-x
.
20x+10y=1300
20x+10(90-x)=1300
20x+900-10x=1300
10x=400
x=40
ANSWER: You spent 40 minutes biking
.
CHECK:
.
y=90-x=90-40=50 . You spent 50 minutes walking.
.
20x+10y=1300
20(40)+10(50)=1300
800+500=1300
1300=1300
.